OpenStudy (anonymous):

suppose you roll a 6 sided die 10 times. What's the probability of getting three fives in those 10 rolls? A. .155 B. .60 C. .30 D. .000618 E. .930 Would you do Binomcdf (10,1/6,3) on your calculator to find this?

OpenStudy (anonymous):

6|10 is.6

OpenStudy (anonymous):

yea i think u would

OpenStudy (dumbcow):

yes you will use binomial distribution, but not the cumulative one. That would tell you prob of getting at least 3 fives in 10 rolls

OpenStudy (dumbcow):

\[\left(\begin{matrix}10 \\ 3\end{matrix}\right) (\frac{1}{6})^3 (\frac{5}{6})^7 = 0.155\]

OpenStudy (anonymous):

okay that makes sense, thank you!